# Volatile-rich Sub-Neptunes as Hydrothermal Worlds: The Case of K2-18 b

### v1

@@ Abstract
Temperate exoplanets between the sizes of Earth and Neptune, known as “sub-Neptunes”, have emerged as intriguing targets for astrobiology. It is unknown whether these planets resemble Earth-like terrestrial worlds with a habitable surface, Neptune-like giant planets with deep atmospheres and no habitable surface, or something exotic in between. Recent JWST transmission spectroscopy observations of the canonical sub-Neptune K2-18 b revealed $\sim$ 1% CH4, $\sim$ 1% CO2, and a non-detection of CO in the atmosphere. While previous studies have proposed that the observed atmospheric composition could help constrain the lower atmosphere conditions and determine the interior structure of sub-Neptunes like K2-18 b, the possible interactions between the atmosphere and a hot, supercritical water ocean at its base remain unexplored. In this work, we investigate whether a global supercritical water ocean, resembling a planetary-scale hydrothermal system, can explain these observations on K2-18 b-like sub-Neptunes through equilibrium aqueous geochemical calculations. We find that the observed atmospheric CH4/CO2 ratio implies a minimum ocean temperature of $\sim 713~{}\rm K$, whereas the corresponding CO/CO2 ratio allows ocean temperatures up to $\sim 1070~{}\rm K$. These results indicate that a global supercritical water ocean on K2-18 b is plausible. While life cannot survive in this ocean, this work represents the first step towards understanding how a global supercritical water ocean may influence observable atmospheric characteristics on volatile-rich sub-Neptunes. Future observations with better constrained NH3 and CO mixing ratios could further help distinguish between possible interior compositions of K2-18 b.
@@ 1 Introduction
Understanding the compositions of exoplanets is crucial for identifying potentially habitable worlds. The majority of confirmed exoplanets are classified as sub-Neptunes – exoplanets between the sizes of Earth and Neptune, that are unlike any planetary body in our solar system (Fressin et al., 2013; Batalha, 2014; Winn & Fabrycky, 2015; Fulton et al., 2017). Because a range of interior compositions can explain the observed masses and radii of sub-Neptunes, it is unknown whether these planets are closer to (1) terrestrial planets with thin atmospheres and potentially habitable surfaces, (2)“mini-Neptunes” which are less massive versions of our solar system’s ice giants, or (3) “in-between state” planets, such as “water worlds” with volatile-rich interiors. Among the sub-Neptune population, a canonical habitable-zone sub-Neptune K2-18 b has captured the attention of the astrobiology community.
@@
K2-18 b has a radius of 2.610 $\pm$ 0.087 $R_{\oplus}$ (Benneke et al., 2019), a mass of 8.63 $\pm$ 1.35 $M_{\oplus}$ (Cloutier et al., 2019), and a bulk density of $2.67^{+0.52}_{-0.47}$ g/cm3 (Benneke et al., 2019) that can be explained by a diverse range of interior structures with varying amounts of H2O and H2/He in a volatile-rich envelope (Madhusudhan et al., 2020; Nixon & Madhusudhan, 2021; Rigby & Madhusudhan, 2024). To discern the most plausible structure, previous studies proposed that the observed atmospheric composition could be key to understanding the conditions at the base of the atmosphere in temperate sub-Neptunes (Hu et al., 2021; Tsai et al., 2021; Yu et al., 2021; Madhusudhan et al., 2023a). Potential atmospheric compositions have been explored for various possible atmosphere-interior structure scenarios (as shown in Figure 1), including a thin-atmosphere rocky world (Tsai et al., 2021; Yu et al., 2021), a thin-atmosphere liquid water world with potential signs of life (known as a “Hycean” world, Hu et al. 2021; Madhusudhan et al. 2023b; Wogan et al. 2024; Tsai et al. 2024), a deep-atmosphere mini-Neptune with no surface (Hu, 2021; Tsai et al., 2021; Yu et al., 2021; Leconte et al., 2024; Wogan et al., 2024), and a deep-atmosphere mini-Neptune with a silicate magma ocean (Shorttle et al., 2024).
@@
Figure 1: Possible conceptual structures for K2-18 b that are explored previously (Hu, 2021; Hu et al., 2021; Tsai et al., 2021; Yu et al., 2021; Madhusudhan et al., 2023b; Shorttle et al., 2024; Wogan et al., 2024) and in this work.
Recent JWST transmission spectroscopy of K2-18 b led to the first detection of CH4 ($\sim$ 1%) and CO2 ($\sim$ 1%) in the atmosphere of a sub-Neptune, along with nondetections of H2O, CO, and NH3 (Madhusudhan et al., 2023b). Despite all of the excellent work that has been done, existing models have struggled to fully explain this observed atmospheric composition. Recent predictions suggest K2-18 b’s interior may be water-rich (Yang & Hu, 2024), and sub-Neptunes with close orbits could host supercritical water oceans (Mousis et al., 2020). Climate models also predict that “Hycean worlds” receiving solar flux with pure H2-He atmospheres will experience runaway greenhouse effects, driving surface water to a supercritical state (Innes et al., 2023; Pierrehumbert, 2023; Leconte et al., 2024). However, the atmospheric interactions with a global supercritical water ocean remain largely unexplored. Water reaches its critical point at 220.64 bar and 647.096 K (Wagner & Pruß, 2002), beyond which there is no distinction between liquid and gas. On K2-18 b, an ocean of supercritical water would manifest as a layer of hot aqueous fluids merging seamlessly with the H2 dominated atmosphere above, potentially creating a global hydrothermal environment. H2 is miscible with supercritical water (Seward & Franck, 1981; Innes et al., 2023; Pierrehumbert, 2023), creating a homogeneous mixture within this transition region. Current photochemical models do not account for interactions between a H2 atmosphere and a supercritical water layer (Hu, 2021; Tsai et al., 2021; Yu et al., 2021; Madhusudhan et al., 2023a, b), highlighting the need for new approaches. Our study addresses this gap by exploring the geochemistry in a supercritical water ocean on K2-18 b to explain the observed abundances of key species in K2-18 b’s atmosphere. We focus on major carbon species as their abundances are well constrained by existing observations (Madhusudhan et al., 2023b).
@@ 2 Methodology
We aim to derive constraints on the boundary conditions at the transition region between the atmosphere and a hypothetical supercritical water ocean. To explore how ocean conditions could drive volatile chemistry, we first calculate the aqueous equilibrium ratios of CH4/CO2 and CO/CO2 in the supercritical water ocean, as detailed in Section 2.1. Given the absence of a distinct boundary between the supercritical water ocean and the gaseous atmosphere, we conduct the aqueous calculations at temperatures and pressures where chemical equilibrium could be achieved. Consequently, the term “ocean temperature” in this study refers to the condition for reaching an apparent chemical equilibrium rather than a physical boundary between the ocean and the atmosphere.
@@ 2.1 Aqueous Chemistry of Carbon
At chemical equilibrium, the molar ratio of CH4/CO2 can be computed as a function of temperature, total pressure, and fluid oxidation state. In an H2-rich environment such as K2-18 b, the oxidation state can be effectively represented by the molality of hydrogen gas dissolved in the ocean. The reaction that describes the chemical equilibrium between CO2 and CH4 in a H2-bearing supercritical fluid can be written as,
@@
where “sc” refers to the supercritical phase, $K$ is the equilibrium constant, $a$ denotes the activity in an aqueous solution, $X$ represents the mole fraction and $m$ is the molality, which describes the number of moles of solute per kilogram of H2O. In Equation 2, following Glein et al. (2008), we make the following approximations: 1) the activity of water is considered equivalent to the mole fraction of water in the supercritical phase; 2) the activity of H2 is approximated by the molality of H2, and 3) the activity ratio of CO2 and CH4 is equivalent to their molality ratio, and thus, their mole fraction ratio (${X}_{\text{CH}_{4}}/{X}_{\text{CO}_{2}}$). Other approximations for water and the aqueous species that reflect thermodynamic standard state conventions can be found in Appendix D.
@@
We chose this reaction because the observed mixing ratios of CH4 and CO2 are the most abundant oxygen and carbon-bearing molecules and also the most well-constrained, while the uncertainty for CO is high. By including CO2 as both a product and reactant, we can express Equation 4 in terms of the carbon species ratios. This choice ensures that our CO/CO2 modeling aligns with the observed CH4/CO2 ratio.
Here we implement the Deep Earth Water (DEW) model, originally developed to study the carbon cycle in Earth’s mantle fluids (Shock & Helgeson, 1988, 1990; Sverjensky et al., 2014), to calculate equilibrium constants under conditions potentially relevant to K2-18 b and determine the molal ratios of CH4/CO2 and CO/CO2 across wide ranges of temperatures, pressures, and fluid oxidation states. Below, we consider “ocean temperatures” ranging from the critical temperature of 374C up to 1000C (647.15 K - 1273.15 K) and three pressure conditions (1 kbar, 5 kbar, and 10 kbar) to model how key geochemical variables affect aqueous ratios involving CO2, CH4, and CO at chemical equilibrium.
@@ 2.2 Constraining the Range of Hydrogen Molality
To model the CH4/CO2 ratios using Equation 2, we require an estimate of the hydrogen molality ($m_{\text{H}_{2}}$, expressed in molal units, $\cdot$(kg H2O)-1), based on the aqueous standard state on K2-18 b. Previous interior structure modeling by Madhusudhan et al. (2020) suggests that the $m_{\text{H}_{2}}$ of K2-18 b varies widely depending on the planet’s bulk composition, from near-zero for water-dominated interiors to higher values for H-He dominated cases. Based on these models, combined with planet formation constraints, we estimate $m_{\text{H}_{2}}$ to plausibly range from 0.15 to 56.55 $\cdot$(kg H2O)-1, although higher values cannot be excluded. For further details on how these limits were derived, refer to Appendix C.
However, the modeling results using $m_{\text{H}_{2}}$ equal and above the upper limit should only be interpreted qualitatively in terms of trends, as higher hydrogen molalities describe increasingly water-poor solutions. These conditions become less applicable to a water-rich supercritical ocean, which is the scenario that can be evaluated using the aqueous chemistry framework of the DEW model. As such, an intermediate hydrogen molality of $m_{\text{H}_{2}}=6.16$ $\cdot$(kg H2O)-1 was also selected to represent a mole fraction of 90% H2O and 10% H2. This value is roughly the upper limit that the DEW model can accommodate while preserving the dilute solution assumptions made in the model.
These hydrogen molalities, $m_{\text{H}_{2}}$ = 0.15, 6.16, and 56.55 $\cdot$(kg H2O)-1, correspond to water mole fractions of $X_{\text{H}_{2}\text{O}}$ = 0.99, 0.90, and 0.42, respectively. For the remainder of the paper, we refer to these compositions as $X_{\text{H}_{2}\text{O}}$ = 99%, 90%, and 42%.
@@ 2.3 Gaseous Carbon Species from Ocean to Atmosphere
While understanding the carbon ratios in the ocean is an important first step, it is also important to understand how those ratios would be translated to the upper atmosphere where atmospheric observations take place. K2-18 b presents a special case due to water condensation, which depletes water vapor in the upper atmosphere and creates a distinct cloud layer of liquid water beneath the observable photosphere. Hence, we must apply Henry’s Law to predict the ratios of atmospheric gases involving CH4, CO2, and CO.
Henry’s Law states that the activity of a solute species in a solvent is proportional to the fugacity of the gas in equilibrium with the solvent at constant temperature and pressure. In ideal fluids, the fugacity is equivalent to the partial pressure. In this context, water serves as the solvent, and gaseous species (CH4, CO2, and CO) dissolve in it. Here, we apply Henry’s Law at the saturation boundary between the liquid and gaseous phases to estimate gas-phase ratios at the base of the cloud layer, where phase partitioning occurs.
@@
Figure 2: (a) Phase diagram of water, showing adiabats for six selected ocean temperatures from 647.15 K to 1273.15 K at a pressure redox equilibrium of 1 kbar. (b) A schematic diagram showing the internal structure of K2-18 b as a Stratified Mini-Neptune, with a supercritical water ocean and a cloud layer in an H2-He dominated atmosphere. Star and circle symbols represent where aqueous and gaseous calculations are performed, respectively.
@@
where ${K}_{\text{H}}$ is the Henry’s Law constant and $p$ is the partial pressure of each species. Using these constants and the aqueous ratios, the gaseous ratio between CH4/CO2 and CO/CO2 can be calculated as:
@@
To determine the temperature and pressure conditions where gas-liquid partitioning occurs, we began by locating the saturation boundary of water using the Reference Fluid Thermodynamic and Transport Properties (REFPROP) program (Huber et al., 2022). Then, we generated adiabatic atmospheric pressure-temperature (P-T) profiles to identify where they intersect the saturation boundary. This process is illustrated in Figure 2a and detailed further in Appendix E. The intersection between the profiles and the liquid-vapor saturation boundary indicates the location of the cloud deck.
@@
Figure 3: Panels (a)-(c) and Panels (d)-(f) represent the calculated aqueous and the gasous CH4/CO2 ratios $X_{\text{H}_{2}\text{O}}$ = 99%, 90%, and 42%. The $X_{\text{H}_{2}\text{O}}$ = 42% results are in dashed lines to imply that results are illustrative, as this concentration describes a hydrogen-rich fluid. The black dashed line and gray region represent the CH4/CO2 ratio and uncertainty presented in Madhusudhan et al. (2023b).

### v2

@@ Abstract
Temperate exoplanets between the sizes of Earth and Neptune, known as “sub-Neptunes”, have emerged as intriguing targets for astrobiology. It is unknown whether these planets resemble Earth-like terrestrial worlds with a habitable surface, Neptune-like giant planets with deep atmospheres and no habitable surface, or something exotic in between. Recent JWST transmission spectroscopy observations of the canonical sub-Neptune, K2-18 b, revealed $\sim$ 1% CH4, $\sim$ 1% CO2, and a non-detection of CO in the atmosphere. While previous studies proposed that the observed atmospheric composition could help constrain the lower atmosphere’s conditions and determine the interior structure of sub-Neptunes like K2-18 b, the possible interactions between the atmosphere and a hot, supercritical water ocean at its base remain unexplored. In this work, we investigate whether a global supercritical water ocean, resembling a planetary-scale hydrothermal system, can explain these observations on K2-18 b-like sub-Neptunes through equilibrium aqueous geochemical calculations. We find that the observed atmospheric CH4/CO2 ratio implies a minimum ocean temperature of $\sim 710~{}\rm K$, whereas the corresponding CO/CO2 ratio allows ocean temperatures up to $\sim 1070~{}\rm K$. These results indicate that a global supercritical water ocean on K2-18 b is plausible. While life cannot survive in such an ocean, this work represents the first step towards understanding how a global supercritical water ocean may influence observable atmospheric characteristics on volatile-rich sub-Neptunes. Future observations with better constrained CO and NH3 mixing ratios could further help distinguish between possible interior compositions of K2-18 b.
@@ 1 Introduction
Understanding the compositions of exoplanets is crucial for identifying potentially habitable worlds. The majority of confirmed exoplanets are classified as sub-Neptunes – exoplanets between the sizes of Earth and Neptune – that are unlike any planetary body in our solar system (Fressin et al., 2013; Batalha, 2014; Winn & Fabrycky, 2015; Fulton et al., 2017). Because a range of interior compositions can explain the observed masses and radii of sub-Neptunes, it is unknown whether these planets are closer to (1) terrestrial planets with thin atmospheres and potentially habitable surfaces, (2)“mini-Neptunes” which are less massive versions of our solar system’s ice giants, or (3) “in-between state” planets, such as “water worlds” with volatile-rich interiors. Among the sub-Neptune population, a canonical habitable-zone sub-Neptune K2-18 b has captured the attention of the astrobiology community (Glein, 2024).
@@
K2-18 b has a radius of 2.610 $\pm$ 0.087 $R_{\oplus}$ (Benneke et al., 2019), a mass of 8.63 $\pm$ 1.35 $M_{\oplus}$ (Cloutier et al., 2019), and a bulk density of $2.67^{+0.52}_{-0.47}$ g/cm3 (Benneke et al., 2019) that can be explained by a diverse range of interior structures with varying amounts of H2O and H2/He in a volatile-rich envelope (Madhusudhan et al., 2020; Nixon & Madhusudhan, 2021; Rigby et al., 2024). To discern the most plausible structure, previous studies proposed that the observed atmospheric composition could be key to understanding the conditions at the base of the atmosphere in temperate sub-Neptunes (Hu et al., 2021; Tsai et al., 2021; Yu et al., 2021; Madhusudhan et al., 2023a). Potential atmospheric compositions have been explored for various possible atmosphere-interior structure scenarios (as shown in Figure 1), including a thin-atmosphere rocky world (Tsai et al., 2021; Yu et al., 2021), a thin-atmosphere liquid water world with potential signs of life (known as a “Hycean” world, Hu et al. 2021; Madhusudhan et al. 2023b; Wogan et al. 2024; Tsai et al. 2024; Cooke & Madhusudhan 2024), a deep-atmosphere mini-Neptune with no surface (Hu, 2021; Tsai et al., 2021; Yu et al., 2021; Wogan et al., 2024), and a deep-atmosphere mini-Neptune with a silicate magma ocean (Shorttle et al., 2024). While the rocky world scenario was found to be inconsistent with observations (Madhusudhan et al., 2020), the water world and mini-Neptune scenarios remain possible.
@@
Figure 1: Possible conceptual structures for K2-18 b that are explored previously (Hu, 2021; Hu et al., 2021; Tsai et al., 2021; Yu et al., 2021; Madhusudhan et al., 2023b; Shorttle et al., 2024; Wogan et al., 2024; Rigby et al., 2024; Cooke & Madhusudhan, 2024) and in this work.
[…]
@@ 2 Methodology
We aim to derive constraints on the boundary conditions at the transition region between the atmosphere and a hypothetical supercritical water ocean. To explore how ocean conditions could drive volatile chemistry, we first calculate the aqueous equilibrium ratios of CH4/CO2 and CO/CO2 in the supercritical water ocean, as detailed in Section 2.1. Given the likely absence of a distinct boundary between the supercritical water ocean and the gaseous atmosphere, we conduct the aqueous calculations at temperatures and pressures where chemical equilibrium could be achieved. Consequently, the term “ocean temperature” in this study refers to the condition for reaching an apparent chemical equilibrium rather than a discrete physical boundary between the ocean and the atmosphere.
@@ 2.1 Aqueous Chemistry of Carbon
At chemical equilibrium, the molar ratio of CH4/CO2 can be computed as a function of temperature, total pressure, and fluid oxidation state. In an H2-rich environment such as K2-18 b, the oxidation state can be effectively represented by the molal concentration of hydrogen gas dissolved in the ocean. The reaction that describes the chemical equilibrium between CO2 and CH4 in an H2-bearing supercritical fluid can be written as,
@@
where “sc” refers to the supercritical fluid phase, $K$ stands for the equilibrium constant, $a$ denotes the activity in an aqueous solution, $X$ represents the mole fraction and $m$ designates the molality, which describes the number of moles of solute per kilogram of H2O. In Equation 2, following Glein et al. (2008), we make the following approximations: 1) the activity of water is considered equivalent to the mole fraction of water in the supercritical phase; 2) the activity of H2 is approximated by the molality of H2, and 3) the activity ratio of CO2 and CH4 is equivalent to their molality ratio, and thus, their mole fraction ratio (${X}_{\text{CH}_{4}}/{X}_{\text{CO}_{2}}$). Other approximations for water and the aqueous species that reflect thermodynamic standard state conventions can be found in Appendix D.
@@
We chose this reaction because the observed mixing ratios of CH4 and CO2 make them the most abundant oxygen and carbon-bearing molecules and they are also the most well-constrained, while the uncertainty on CO abundance is high. By including CO2 as both a product and reactant, we can express Equation 4 in terms of the carbon species ratios. This choice ensures that our CO/CO2 modeling always aligns with the observed CH4/CO2 ratio.
Here we implement the Deep Earth Water (DEW) model, originally developed to study the carbon cycle in Earth’s mantle fluids (Shock & Helgeson, 1988, 1990; Sverjensky et al., 2014), to calculate equilibrium constants under conditions potentially relevant to K2-18 b and determine the molal ratios of CH4/CO2 and CO/CO2 across wide ranges of temperatures, pressures, and fluid oxidation states. Below, we consider “ocean temperatures” ranging from 374C up to 1000C (647.15 K - 1273.15 K), where we reference 647.15 K as the critical temperature of pure water, and three pressure conditions (1 kbar, 5 kbar, and 10 kbar) to understand how key geochemical variables affect aqueous ratios involving CO2, CH4, and CO at chemical equilibrium.
@@ 2.2 Constraining the Range of Hydrogen Molality
To model the CH4/CO2 ratios using Equation 2, we require an estimate of the hydrogen molality ($m_{\text{H}_{2}}$, expressed in molal units, $\cdot$(kg H2O)-1), based on the aqueous standard state on K2-18 b. Previous interior structure modeling by Madhusudhan et al. (2020) suggests that the overall $m_{\text{H}_{2}}$ of K2-18 b varies widely depending on the planet’s bulk composition, from near-zero for water-dominated interiors to higher values for H-He dominated cases. Based on these models, combined with planet formation constraints, we estimate $m_{\text{H}_{2}}$ to plausibly range from 0.14 to 56.55 $\cdot$(kg H2O)-1, although higher values cannot be excluded. For further details on how these limits were derived, refer to Appendix C.
However, the modeling results using $m_{\text{H}_{2}}$ equal and above the upper limit should only be interpreted qualitatively in terms of trends, as higher hydrogen molalities describe increasingly water-poor solutions. These conditions become less applicable to a water-rich supercritical ocean, which is the scenario that can be evaluated based on the assumptions inherent to our aqueous chemistry framework (See Appendix D). As such, an intermediate hydrogen molality of $m_{\text{H}_{2}}=6.16$ $\cdot$(kg H2O)-1 was also selected to represent a mole fraction of 90.0% H2O and 10% H2, excluding He in this case. This value is roughly the upper limit that the DEW model can accommodate while preserving the dilute solution assumptions made in the model.
These hydrogen molalities, $m_{\text{H}_{2}}$ = 0.14, 6.16, and 56.55 $\cdot$(kg H2O)-1, correspond to hydrogen mole fractions of $X_{\text{H}_{2}}$ = 0.002, 0.100, and 0.433 and water mole fractions of $X_{\text{H}_{2}\text{O}}$ = 0.996, 0.900, and 0.424, respectively. For the remainder of this paper, we refer to these compositions as $X_{\text{H}_{2}\text{O}}$ = 99.6%, 90.0%, and 42.4%.
@@ 2.3 Gaseous Carbon Species from Ocean to Atmosphere
While understanding the carbon ratios in the ocean is an important first step, it is also important to understand how those ratios would be translated to the upper atmosphere where atmospheric observations are made. K2-18 b presents a special case due to water condensation, which depletes water vapor in the upper atmosphere and creates a distinct cloud layer of liquid water beneath the observable photosphere (Benneke et al., 2019; Madhusudhan et al., 2023a). Hence, we must apply Henry’s Law to predict the ratios of atmospheric gases involving CH4, CO2, and CO.
Henry’s Law states that the activity of a solute species in a solvent is proportional to the fugacity of the gas in equilibrium with the solution at constant temperature and pressure. In ideal fluids, the fugacity is equivalent to the partial pressure. In this context, water serves as the solvent and gaseous species (CH4, CO2, and CO) dissolve in it. Here, we apply Henry’s Law at the saturation boundary between the liquid and gaseous phases to estimate gas-phase ratios at the base of the cloud layer, where phase partitioning occurs.
@@
Figure 2: (a) Phase diagram of pure water, showing adiabats for six selected ocean temperatures from 647.15 K to 1273.15 K at a pressure of redox equilibrium of 1 kbar as an example. This plot is most relevant to cases with lower H2 abundances in the fluid envelope. Liquid water will be saturated at lower temperatures for the highest H2 abundance considered in this work (see Appendix). (b) A schematic diagram showing the upper interior structure of K2-18 b with a miscible supercritical H2+H2O layer (Benneke et al., 2024), or (c) with phase-separated H2-rich and H2O-rich layers and a cloud layer in an H2-He dominated atmosphere. Star and circle symbols represent where aqueous and gaseous calculations are performed, respectively.
@@
where ${K}_{\text{H}}$ represents the Henry’s Law constant and $p$ corresponds to the partial pressure of each species in bars. Using these constants and the aqueous ratios, the gaseous ratio between CH4/CO2 and CO/CO2 can be calculated as:
@@
To determine the temperature and pressure conditions where gas-liquid partitioning occurs, we began by locating the saturation boundary of water using the Reference Fluid Thermodynamic and Transport Properties (REFPROP) program (Huber et al., 2022). Then, we generated adiabatic atmospheric pressure-temperature (P-T) profiles to identify where they intersect the saturation boundary. This process is illustrated in Figure 2a and detailed further in Appendix E. The intersection between the profiles and the liquid-vapor saturation boundary indicates the approximate location of the cloud deck.
@@
Figure 3: Panels (a)-(c) and panels (d)-(f) show the calculated aqueous and the gaseous CH4/CO2 ratios at $X_{\text{H}_{2}\text{O}}$ = 99.6%, 90.0%, and 42.4%. The $X_{\text{H}_{2}\text{O}}$ = 42.4% results are in dashed lines to imply that results are illustrative, as this concentration describes a hydrogen-rich fluid (see Appendix D). The black dashed line and gray region represent the CH4/CO2 ratio and uncertainty presented in Madhusudhan et al. (2023b).